a.Define marginal efficiency of capital.How does it determine the volume of investment?
b.Given the saving function S=-1.0 +0.2Y and autonomous investment I=$500 million .
1.Find the level of consumption .
2.If investment decreases permanently by $50 million.What is the new level of consumption?
1
Expert's answer
2015-01-15T09:22:57-0500
a. The marginal efficiency of capital (MEC) is the net rate of return that is expected from the purchase of additional capital. It is calculated as the profit that a firm is expected to earn considering the cost of inputs and the depreciation of capital. It is influenced by expectations about future input costs and demand. The MEC and capital outlays are the elements that a firm takes into account when deciding about an investment project.
b. If S=-1.0 +0.2Y and autonomous investment I=$500 million, then:
1. The level of consumption is:
C = c*Y, c = 1 - s = 1 - 0.2 = 0.8
I = S
-1 + 0.2Y = 500
Y = $2505 million
C = 0.8*2505 = $2004 million.
2.If investment decreases permanently by $50 million. The new level of consumption will be:
C = 0.8*2280 = $1824 million